3. Given an urn containing 2 green, 3 white, and 4 red balls. If two balls are drawn and the first ball is not replaced before the second is drawn, what is the probability that the first ball is white and the second is red?

a. 2/3 b. 7/17 c. 4/27 d. 1/6 e. 2/9
Answer #1

The probability that the first ball is white and the second is red = 3/9*4/8 = 1/6